Assume that equilibrium real GDP is $ 800 billion, potential real GDP is $ 950 billion, the MPC is .80, and the MPI is .40.

a. How much taxes fall to eliminate the GDP gap?

b.If government spending and taxes both change by the same amount, how much must they change to eliminate the recessionary gap?
Answer #1

a. GDP gap = 950 - 800= 150 billion

Tax multiplier = 0.8/(1-0.8 + 0.4) = 1.33

Taxes need to decrease by 150/1.33= 112.5

b. Balanced budget multiplier is 1.

Government spending and taxes should change by $150 billion